OpenStudy (anonymous):

Please Help! Part A: Write an equation in two variables in the standard form that can be used to describe the velocity of a car at different times. Part B: How can you graph the equation obtained in Part A for the first two hours?

OpenStudy (ranga):

We will have to assume that the acceleration is constant to answer this problem. If the acceleration is constant the velocity will be a straight line and the equation for velocity will be: v(t) = at + b where a and b are unknown constants. But we know when t = 2, v = 50 and when t = 6, v = 54. Substitute v and t in v = at + b and solve for a and b. Part B: Plot the graph. Find v(t) for t = 1, 2, 3, 4, 5, 6 and 7 and plot the points.

OpenStudy (anonymous):

Thank you! And is v(t) = at + b standard form?